Please help. Solve. x^2 + 16x + 69 = 6
Answer by stanbon(75887)

Solve. x^2 + 16x + 69 = 6
--------
x^2 + 16x + 63 = 0
(x+9)(x+7) = 0
x = -9 or x = -7
========================
Cheers,
Stan H.